### v1

# Graph-Based Captioning: Enhancing Visual Descriptions by Interconnecting Region Captions

Abstract
Humans describe complex scenes with compositionality, using simple text descriptions enriched with links and relationships. While vision-language research has aimed to develop models with compositional understanding capabilities, this is not reflected yet in existing datasets which, for the most part, still use plain text to describe images. In this work, we propose a new annotation strategy, graph-based captioning (GBC) that describes an image using a labelled graph structure, with nodes of various types. The nodes in GBC are created using, in a first stage, object detection and dense captioning tools nested recursively to uncover and describe entity nodes, further linked together in a second stage by highlighting, using new types of nodes, compositions and relations among entities. Since all GBC nodes hold plain text descriptions, GBC retains the flexibility found in natural language, but can also encode hierarchical information in its edges. We demonstrate that GBC can be produced automatically, using off-the-shelf multimodal LLMs and open-vocabulary detection models, by building a new dataset, GBC10M, gathering GBC annotations for about 10M images of the CC12M dataset. We use GBC10M to showcase the wealth of node captions uncovered by GBC, as measured with CLIP training. We show that using GBC nodes’ annotations—notably those stored in composition and relation nodes—results in significant performance boost on downstream models when compared to other dataset formats. To further explore the opportunities provided by GBC, we also propose a new attention mechanism that can leverage the entire GBC graph, with encouraging experimental results that show the extra benefits of incorporating the graph structure. Our datasets are released at https://huggingface.co/graph-based-captions.
Introduction
The availability of huge paired image/caption datasets has revolutionized our ability to produce joint vision-language embedding, paving the way for tasks like efficient caption-guided image generation within powerful multimodal foundation models [50, 36, 2, 38]. The quality and granularity of these datasets plays, therefore, a crucial role. While quality can be addressed by filtering out data [53, 17, 22] or, inversely, by improving caption quality through recaptioning [35, 14, 45, 16], there is ample interest in the community to provide more detailed, fine-grained information for each image [42, 3, 7, 15]. To obtain better annotations, we draw inspiration from compositionality, a fundamental characteristic of human perception that is reflected in the natural language used to describe our surroundings [12, 20, 30, 29, 4, 10]. Compositionality plays an especially important role when examining larger images found in the wild, which have a rich coarse-to-fine, hierarchical structure, commonly represented as a scene graph [31]. While scene graphs have been successfully applied to image retrieval [54, 31], generation [44, 18], and pre-training [26, 28], the scale of scene graph dataset is typically small. For instance, Visual Genome [33] only contains around k images.
Contributions.
To overcome the limitations of existing datasets and annotation formats that either struggle to represent the hierarchical nature of scenes or are of small size and lack flexibility in their description, this paper makes a series of contributions as summarized below.
1. We propose graph-based captioning (GBC), a new vision-language data format that captions images with a graph-based structure akin to scene graphs while retaining the flexibility and intuitiveness of plain text description. GBC contains four types of nodes: (1) an image node with captions of the entire image, (2) entity nodes that contain descriptions of individual objects, (3) composition nodes that link objects in the images of the same type, and (4) relation nodes that describe the spatial (“the tree is to the left of the tower”) or semantic (“The branch is covered in snow”) relationships between objects of different types (§ 3.1).
2. We design a workflow to produce GBC annotations at scale. Inspired by recent recaptioning approaches [35, 14, 45, 16], we generate GBC annotations using the OSS LLaVA-1.6 [42, 43]. First, LLaVA generates short and long captions for the entire image, used to extract entities. Then, an object detection model (YOLO-World [9]) is employed to find bounding boxes for each entity. Subsequently, the same procedure is recursively invoked to produce a GBC for each proposal. Finally, LLaVA-1.6 is prompted to produce composition and relation captions that connect multiple entity nodes (§ 3.2).
3. We create large-scale GBC dataset containing million images with words per image using that workflow. While ours is the first vision-language dataset that contains structured captions, a few recent datasets contain dense annotations, and only [65] has a scale that is similar to ours. Drawing on graph-encoders, we also design a baseline architecture to utilize this structure, using a new attention mechanism, SAHA (SAHA) (§ 4).
4. We demonstrate experimentally in § 5 that the diversity of captions found in GBC nodes improves CLIP model performance across image-to-text retrieval, text-to-image retrieval, compositionality, and semantic segmentation tasks, while retaining comparative performance on zero-shot ImageNet classification. Our ablations highlight that it is not the density or scale of our dataset, but the structured nature itself, that leads to better performance. Remarkably, we observe that composition and relation nodes, which can only be obtained through the GBC workflow, boost performance. Finally, we perform ablation on the influence of annotation format on retrieval performance using a set-aside test set from GBC. In this case, we see that SAHA, when fed GBC annotations, provides comparable or even better performance than that obtained when describing an image with detailed captions, suggesting that GBC can be a promising alternative to traditional image captioning formats.
Related works
In this section, we discuss related works on vision-language datasets. We refer the readers to Appendix A for works that are specific to CLIP [50] training.
Vision-language datasets. First vision-language datasets were manually built using human annotations, such as Flickr30k [69], COCO [41] and Visual Genome [33]. This yielded annotations of high quality, but unfortunately of short length, and in limited amounts (with no dataset containing more than 130k images). Several studies have then demonstrated the benefits of using larger scale datasets obtained by crawling the web, such as YFCC100M [59], RedCaps [13], or Wikipedia-based image-text dataset (WIT) [57]. The quality of these data became a concern when it was noticed that in some situations the caption was only loosely related (or not related at all) with the image, which can be detrimental to the overall performance [52]. This motivated researchers to use automatic filtering procedures to select higher-quality data samples, like in Localized Narratives [48] or Conceptual Captions (CC3M) [56], and its successor CC12M [6]. These efforts have reached billion scale with LAION-5B [53], and LAION-CAT [49]. In a similar vein, Meta-CLIP [68] reproduces the processing of the seminal CLIP paper [50] on a subset of the Common Crawl dataset, SemDeDup [1] relies on the embeddings provided by foundation models to filter data and remove duplicates, while DFN [17] uses filtering networks trained on high quality data to extract subsets of Common Crawl.
VL datasets with dense captioning. It was noticed recently that using entirely generated captions from raw images, as in DAC [14] and AS-1B [65], could improve results over filtering approaches. These datasets are characterized by their long and detailed captions that describe every element within a scene. Complementing these efforts, Urbanek et al. [61] introduced DCI, a dataset featuring similarly dense annotations but curated by humans and on a smaller scale. Alternatively, DOCCI [46] focuses on a set of only 15k high quality, high resolution, paired image-captions, manually selected and annotated by one of the authors, with typical caption length of more than words. In the ImageInWords [21] dataset, captions are iteratively improved by humans, on top of previously human or machine annotated captions, yielding 9K densely captioned images.
Improving image annotations with graph-based captioning
We introduce in this section our new captioning format to represent an image, explain how we can use any off-the-shelf MLLM (MLLM) and open-vocabulary detection model to obtain such captions, and briefly describe the two datasets GBC1M, and GBC10M that we construct following the proposed workflow. Additional details about the data preparation process and the datasets can be found in the Appendices B and C.
3.1 Representing an image with graph-based captions
To encode the structured information contained in an image, we propose to represent each image as a DAG (DAG), denoted as . Each node of the graph is associated with a bounding box. Starting with the root node, which corresponds to the entire image (image node), other nodes can either hold a set of objects (composition node and relation node), or a single object in the image (entity node). Moreover, to benefit from the expressive power of natural language descriptions and to ensure smooth integration of our annotations into the existing ecosystems of methods that rely primarily on image-text pairs, we label each node with a set of captions .
The edges, on the other hand, are used to encode the hierarchy between the nodes. More specifically, there is an edge from to only if the content associated to is part of the content associated to . This relation is also reflected by the edge label which should appear in the captions of the source node and be able to represent the object(s) associated to the target node.111Ideally, we would also like to distinguish between multiple appearances of the same text in a caption. However, this is not explicitly handled by our current dataset construction workflow so we omit it here.
An examplar GBC, generated automatically through our workflow is provided in Figure 1. It should be noted that the only manual addition in that graph comes from the "title label" of each node, which is obtained by taking the union of labels found in its incoming edges. Compared to the standard scene graph annotation, the use of node captions provides flexibility to describe complex concepts, while the underlying graph still captures the inherent structure of the image. Our dataset, whose construction is detailed in Section 3.2 next, includes several different types of captions tailored to the structure of the DAG. At the root image node, we provide both detailed and short captions to cater to varying levels of granularity. Captions at composition nodes and relation nodes explicitly describe the arrangement and interaction of multiple objects, while the captions at the entity nodes provide detailed description of a single object.
3.2 GBC dataset construction workflow
We show how to produce GBC annotations automatically, using any pre-trained MLLM and open-vocabulary detection model. This results in a workflow that is comparable, in compute time and complexity, to that of other widespread recaptioning approaches. At a high level, we use a MLLM model to provide captions and identify potential entity nodes, followed by a detection model to provide bounding box coordinates for these entities.
Data annotation. Our overall process to annotate a single image is shown in Figure 2. To account for the different types of nodes, we design four query templates as listed below:
Image query: We ask the model to provide detailed caption for the image, identify prominent elements, and summarize the long caption with a concise one that contains all these elements. The identified elements are then passed to the detection model to obtain the bounding boxes.
Entity query: For each bounding box, we crop out the region and ask the model whether a specific object appears in the cropped image. Moreover, we also ask the model to describe the object and identify prominent elements of the object when it is present. The identified elements are again passed to detection models for detection.
Composition query: In the case where multiple bounding boxes are returned for a single type of object, we ask the model to describe the composition of these objects with an annotated image.
Relation query: For image or entity nodes with more than two children, we ask the model to describe the relations between its children.
Provided that there is no guarantee that all the detected objects would end up as a node in the graph—consider the case where the MLLM says that the object is not present or just fails to reply in the correct format—we split the entire process into two stages, and we only perform composition queries and relation queries after discovering all the entity nodes. Finally, to improve efficiency and to reduce redundant information, we train two dedicated classifier on top of Jina Embeddings [23] to decide whether a piece of text is suitable for object detection and whether two texts can represent the same object in an image. The former is applied to every identified element while the later results in merging of nodes when a new query targets a region that has already been queried with similar texts.
3.3 GBC1M and GBC10M
| GBC1M | GBC10M | |
| # Images | 1,013,592 | 10,138,757 |
| # Vertices / Image | 12.12 | 12.24 |
| # Edges / Image | 22.28 | 21.81 |
| # Captions / Image | 17.40 | 17.67 |
| # Words / Image | 593.14 | 533.98 |
| Average Graph Diameter | 4.55 | 4.41 |
Following the process outlined in Section 3.2, we annotate the CC12M dataset [6] with graph-based captions using LLaVA-1.6 [42, 43] as the MLLM and Yolo-World [9] as the open-vocabulary detection model. Specifically, we construct two sets of annotations: GBC1M for a subset of around 1M of images, with all the queries performed with the Yi-34B version of LLaVA-1.6, and GBC10M for a subset of around 10M of images, with LLaVA-1.6 Yi-34B for image and composition queries, and LLaVA-1.6 Mistral-7B for entity and relation queries.222Our larger dataset does not cover the entire CC12M both because some images were no longer accessible at the time we accessed the images, and because we discard images for which the MLLM model’s reply to the image query does not comply with the prescribed format.
We provide statistics of the above two datasets in Table 1. We note that these two datasets have very similar per-image statistics, with the number of words being the only exception, as LLaVA-1.6 Yi-34B tends to provide longer descriptions than LLaVA-1.6 Mistral-7B. Moreover, our datasets use an average number of around 500 words to describe each image. This is comparable to other dataset with rich annotations such as DCI (1111 words/img) [61] and DOCCI (136 words/img) [46]. We also compute the CLIP scores between the captions and their corresponding regions using the DFN-5B CLIP model [17], and we report their distribution for the GBC10M dataset in Figure 3. We note that the original CC12M caption achieves the highest CLIP scores, followed by the short synthetic caption for the entire image. This can be explained by the fact that in these two cases, the involved image-caption pairs more closely align with the training data of standard CLIP models.
Encoding GBC via structure-aware hierarchical attention
Alongside many ways to leverage GBC annotations, as we shall present in Section 5, we propose a simple text encoder architecture to incorporate structural information encoded in GBC graph along with node captions. Specifically, we present SAHA (SAHA) block which treats each caption as an individual sample, and introduces an additional cross-attention layer that enforces the captions to attend to their children.
Formally, we consider a caption graph with vertices and edges such that if and only if , , , and the label is included within the caption . In words, each vertex in the graph represents a caption from a node of the original graph and there is an edge from one caption to another only if the second caption describes part of the first caption. After tokenization of the captions, we can map the edge labels to a set of token positions of the source caption, which we write as . Then, the target caption annotates the source caption via the tokens at positions . Therefore, we can simply consider cross-attention with queries from these tokens and keys and values from the target caption. We illustrate this idea in Figure 4, where we zoom in on the additional cross-attention layer (SACA) on the right side of the figure. For completeness, we also provide the corresponding mathematical formula in Appendix D.
Our text encoder then stacks a number of SAHA blocks, effectively interleaving the vanilla self-attention layers that process, local, intra-caption information, with the SACA (SACA) layers that process global, inter-caption information in a structure-aware manner. Furthermore, the model acts as a classic text encoder in the absence of edges in the graph, i.e., when , dropping all SACA blocks.
Complexity analysis. To estimate the computational complexity of our approach, we assume that the captions have a fixed sequence length . Then, implementing SACA using masked cross-attentions between captions leads to a total complexity of . Additionally, we must account for the self-attention operations, resulting in a combined complexity of . Provided that many of the graphs in our dataset are trees, we have and the complexity simplifies to . In contrast, a naive approach that performs self-attention on the concatenated set of all captions would result in a significantly higher complexity of .
Experiments
We present in this section a comprehensive set of experiments to benchmark different image annotation schemes. Specifically, using CLIP model training as the main task, we show that GBC annotations can bring improvements on a range of benchmarks across classification, retrieval, and dense prediction tasks, compared to existing annotation schemes (Section 5.3). On retrieval tasks, we demonstrate how GBC allows one to encode denser, more descriptive textual information to better represent images as shown by the performance gain compared to existing annotation formats (Section 5.4). Missing experimental details are provided in Appendix E.
5.1 Annotation formats
We outline below the different types of image annotations that are considered in our experiments, each providing different opportunities to leverage information from the image.
Short caption. Each image is paired with a short caption, as in common image-text datasets.
Long caption. One can improve image description using a longer caption. We use long captions in our dataset. They are of 110 words on average, as compared to short captions, of only 28 words on average. We extend the context length of text encoders in CLIP models from 77 to 512 for this setup.
Region captions. Alternatively, more captions can be provided for an image, especially those that describe a specific region of the image. While this format includes all region captions, it does not include the relational information between region descriptions found in GBC.
Graph-based captions. Finally, we consider the GBC format as proposed in Section 3.1. The GBC format includes region captions, but also provides additional information, stored in relation and composition nodes. With this in mind, we explore three different ways to leverage GBC annotations:
A direct way to leverage GBC is to treat captions for all nodes in the graph as positive texts for the image, i.e. as what we do for region captions. We refer to such method as GBC-captions.
Another strategy is to traverse from the root image node through the graph and concatenate the captions at the visited vertices into a single long caption. We then train a CLIP model with 512 context length in the standard fashion. We refer to this method as GBC-concat.
To fully benefit from the graph information, when available, we leverage the SAHA block proposed in Section 4. This allows us to encode the entire graph into text embeddings that also contain information from their respective subgraph. We refer to this method as GBC-graph.
As GBC annotation encapsulates all existing short, long, and region caption formats, we are able to instantiate all the setups by using only a subset of annotation available in our curated GMC10M dataset. Specifically, taking only the short or detailed caption at the root image node creates the short and long caption setup, respectively. To mimic the region caption setup, we drop the relation and composition captions from GBC annotations. By turning GBC into these settings, we ensure the same quality of text annotations across different methods.
l|cc|cc|c|c|c[colortbl-like]
Annotation
Flickr-1k
MSCOCO-5k
ImageNet
SugarCrepe
ADE20K
T2I
I2T
T2I
I2T
CC12M 46.4 64.6 25.0 39.4 39.2 72.9 41.7
Short 56.3 73.2 30.7 46.7 38.8 76.0 42.0
Long 56.4 75.2 31.8 50.1 39.6 77.0 42.8
Region 58.3 76.6 31.5 49.1 38.5 75.6 43.5
GBC-captions 60.6 79.3 34.1 51.9 40.8 76.7 45.0
GBC-concat 56.1 76.0 31.4 48.5 39.0 75.7 42.1
GBC-graph 58.1 76.2 32.2 49.5 38.2 75.2 43.8
5.2 Experimental setup
We perform CLIP training on our GBC10M dataset, while leaving out 10,151 samples as the test set. Following common practice, we use the CLIP score computed by a pre-trained CLIP model [17] to filter our training set, discarding the 5% of captions with the lowest scores for each type. In addition, we retain the original CC12M captions associated with each image. Specifically, in all setups, both the original caption and the short synthetic caption are consistently used as positive texts for the image during training. This prevents the severe distribution shifts that could occur from using only long or region captions when evaluating on standard benchmarks.
Objective. To pair an image with multiple captions in training CLIP models, we adopt a multiple-positive contrastive loss in the spirit of LaCLIP [16] and DAC [14]. Briefly speaking, compared to standard CLIP objective, the multiple-positive loss sums over the loss on each positive captions of an image while all the captions from the images in the same batch are used in the normalization term.
Model and hyperparmeters. We use the standard CLIP ViT-B/16 model, with the only difference of longer context length of text encoder for long caption and GBC-concat, and a replacement of the vanilla transformer block by the SAHA block in text encoder for GBC-graph. We fix the global batch size (i.e., number of images in each batch) to 4,096 for all the methods. The models are trained for 45,000 steps with AdamW and cosine scheduler at a learning rate of . This roughly correspond to 20 epochs of training. We evaluate at the EMA checkpoint at epoch 10, as we observe that further training provides little to no improvement in performance across the benchmarks.
5.3 Evaluations on existing benchmarks
We compare the CLIP models derived from different annotation schemes on an array of evaluation benchmarks, including: Flickr-1k [47] and MSCOCO-5k [41] for zero-shot retrieval, ImageNet [51] for zero-shot classification, SugarCrepe [27] for compositional understanding evaluation, and ADE20k [76] for semantic segmentation that measures models’ dense prediction performances. Table 5.1 illustrates our results, from which we draw the following two key insights.
GBC annotation leads to clear performance gains by encoding relational information. Table 5.1 demonstrates that training models with more detailed textual information, such as long captions or region captions, consistently enhances downstream performance, particularly in retrieval tasks and dense prediction. However, the most significant improvements are seen with GBC-captions, which augment traditional region captions with relational and compositional descriptions. Given that the GBC workflow is uniquely positioned to provide these, this result demonstrates the soundness of GBC, capturing valuable insights not present in conventional captions.
How the captions are used matters. Compared to GBC-captions, the improvements achieved by GBC-concat and GBC-graph on these benchmarks are of a smaller margin. This indicates that the way GBC annotations is used significantly impacts performance. Specifically, this worse performance is likely due to a mismatch between training and evaluation. For instance, the graph information that would benefit GBC-graph most is not provided in any of these benchmarks. We address this discrepancy below.
5.4 Evaluation on GBC test set
l|cc|cc|cc|cc|cc[colortbl-like]
Annotation
Short
Long
GBC-captions
GBC-concat
GBC-graph
T2I
I2T
T2I
I2T
T2I
I2T
T2I
I2T
T2I
I2T
Short 85.8 86.2 85.0 87.2 57.3 37.0 87.4 88.2 - -
Long 86.4 87.5 95.4 95.7 44.3 33.1 90.5 91.4 - -
Region 85.3 86.1 85.5 88.2 91.5 79.3 89.5 90.0 - -
GBC-captions 86.8 87.6 87.2 89.6 91.3 80.9 90.1 91.0 - -
GBC-concat 86.1 86.5 92.7 93.5 57.5 37.9 94.6 94.9 - -
GBC-graph 84.3 85.2 84.9 87.6 90.8 79.3 89.1 89.7 95.7 96.1
To assess the effectiveness of different annotation formats, we provide the model with these annotations at test time. Annotations that better describe the images should, ideally, result in better retrieval performance when they are used. Specifically, we use our own test set and consider performing retrieval with the various types of annotations presented in Section 5.1. Note, however, that when using region captions or GBC-captions, no single text embedding can naturally encompass all the relevant information. To address this limitation, we perform retrieval based on the average CLIP score between the image embedding and the text embeddings of the provided captions in this setup. We report our results in Section 5.4, where the rows correspond to the annotations at training time and the columns correspond to the annotation at test time. Unsurprisingly, we see a strong tendency that when a model is trained with a certain annotation format, it performs the best when we use the same format for retrieval. Among the few exceptions, we note that models trained to pair with shorter captions may have better performance when concatenation of short captions is provided at test time. This leads us to the following two observations.
Denser textual information improves retrieval performances. The table clearly shows that training with richer annotations—such as long captions, GBC-concat, or GBC-graph—enhances retrieval performance. This improvement suggests that these methods provide a more effective representation of the images. Specifically, GBC-graph yields the best performance, indicating that the proposed GBC format consists in a viable alternative to the commonly used detailed captions.
Simple augmentation during training does not allow to exploit additional information when available. Our observations from Section 5.3 show that treating all captions as independent positives yields the best performance on existing benchmarks. However, there is no evidence that this method could harness the richer information from multiple captions when they are provided together in test time. Indeed, whether we use average CLIP score or concatenation, the retrieval performance of these methods significantly lags behind those methods that are trained directly with captions that individually encompass rich information.
5.5 Ablation studies
In addition to our main experiments, we have conducted extensive ablation studies on both the training and evaluation of our models. We present two of them here and defer the remaining to Appendix F.
Retrieval with multiple captions using maximum CLIP score. An alternative to the mean CLIP score we considered in Section 5.4 is to take the maximum, for which we report the results in Section 5.5. Compared to taking the average, using the maximum is more robust to low CLIP scores, and thus gives better results when the model is not trained to match the image with all local captions, as with Short, Long caption, and GBC-concat. Nonetheless, despite these differences, the overall retrieval performance still significantly lags behind that achieved using a single caption.
l|cc[colortbl-like] Annotation T2I I2T
Short 71.5 78.1
Long 72.9 80.1
GBC-captions 86.4 87.3
GBC-concat 73.3 79.7
GBC-graph 85.1 85.7
{NiceTabular}
l|cc|cc|cc[colortbl-like]
Annotation
DCI-Long
DCI-concat
ShareGPT4V-15k
T2I
I2T
T2I
I2T
T2I
I2T
Long 53.6 53.3 63.3 65.8 93.4 93.9
GBC-captions 42.5 43.3 64.3 63.8 78.7 82.1
GBC-concat 51.4 52.3 69.0 70.8 89.5 91.4
Table 5: Image and text retrieval performance on DCI and a 15k subset of ShareGPT4V-cap100k of our models trained on longer captions. We also include GBC-captions that by design can only handle short captions as a baseline for comparison.
Retrieval with long captions. To complement the results presented in Section 5.1,
we evaluate the retrieval performance of our extended context models on datasets with dense annotations.
We focus specifically on ShareGPT4V [8], which offers GPT-style detailed captions closely resembling those obtained from LLaVA, and DCI [61], containing human-annotated detailed and region captions.
The latter allows us to perform retrieval using either detailed captions or concatenated short captions, as we did in Section 5.4.
Our results shown in Section 5.5 demonstrates that the close caption distribution with ShareGPT4V effectively enables strong retrieval results for models trained on our long captions.
However, potentially due to the distribution shift, all the models perform badly on DCI retrieval with long captions.
In this setup, using concatenated captions for training and retrieval significantly outperformed other baselines, indicating the broader benefit of the concatenation approach.
Conclusion
We propose graph-based captioning (GBC) as a new image-text annotation format, and curated GBC1M and GBC10M datasets.
Grounding on CLIP model training, we propose various baseline methods to utilize the GBC datasets.
Via our experiments, we demonstrate that training with GBC leads to improvements in various benchmarks compared to models derived from traditional annotation formats.
This suggests that GBC provides richer textual information than existing annotation schemes
and
could hence be a valuable foundation for developing more advanced vision-langauge models across various applications.
Acknowledgements
The authors are thankful to Alaa El-Nouby, Enrico Fini, Rick Chang, Shuangfei Zhai, Jiatao Gu, Yizhe Zhang, Hanlin Goh, Josh Susskind, and Vaishaal Shankar for the valuable feedback that have been provided throughout the course of the project.
Appendix A Related works, limitations, and societal impact
This appendix delves deeper into the broader context of our study, examines additional related works, discusses the limitations of our methodologies, and explores its potential societal impacts.
A.1 Additional related works
In this section, we include works related to CLIP training.
CLIP with recaptioning. CLIP [50] is a seminal vision-language model that utilizes text and image encoders to generate joint latent representations.
While there is an extensive body of literature on CLIP training—ranging from modifications in the objective [70, 37], data augmentation techniques [40, 19], to training procedures [72, 58]—it is impossible to cover all developments comprehensively here.
Among these, particularly relevant to our work is the recent trend that highlights the benefits of enhancing caption quality through dedicated models.
For instance, VeCLIP [35] enriches image alt-text with outputs from LLaVA, while similar recaptioning strategies have also been explored by Nguyen et al. [45], Doveh et al. [14], and Vasu et al. [64].
On the other hand, LaCLIP [16] employs LLaMA [60] to rewrite captions.
Going further, SynthCLIP [24] leverages a dataset with entirely generated captions and images for CLIP training.
CLIP with additional annotations. There has been a plethora of research on training CLIP models with diverse annotations such as long captions, region captions, and scene graphs.
As for long captions, DreamLIP [74] proposes to sample sub-captions from the long description to construct multiple positive pairs, while Long-CLIP [73] addresses CLIP’s 77-token limitation by modifying the positional encoding to accommodate longer text sequences during fine-tuning.
Meanwhile, region annotations with varying granularity have been considered by works including GLIP [39], X-VLM [71], and RegionCLIP [75].
Their objectives match features of image crops to their specific descriptions.
Efforts that aim to improve CLIP training with the help of scene graphs include CLIP-SGVL [26] and Structure-CLIP [28].
The former integrates scene graphs to define additional objective for image encoder, while the later uses scene graphs to guide the generation of negative captions, and to enrich the text encoder with additional contextual information.
A.2 Limitations and perspectives
We discuss below the limitations of our works from three different perspectives, the procedure and format, the datasets, and the experiments.
These limitations also naturally point to several future directions that are to be explored.
A.2.1 Limitation concerning the GBC procedure and format
While GBC remains a versatile high-level annotation format that in principle applies to any image, its design is inherently tied to the coarse-to-fine and compositional nature of natural images.
This design orientation means that GBC is not necessarily the most suitable for certain types of images such as scientific imagery, homogeneous patterns, or abstract art.
Specifically, scientific imagery often requires annotations that convey precise, quantifiable data rather than relational or descriptive text.
This limitation highlights the need for tailored approaches to different visual content categories to address their unique characteristics.
A.2.2 Limitation concerning the GBC datasets
Our datasets are curated with the help of LLaVA and Yolo-World, and hence inherit their limitations.
This includes but is not limited to, the bias and hallucination from LLaVA captioning, incorrectly identified objects from Yolo-World, and the inability of Yolo-World to recognize certain object category (see Section C.3 for concrete examples).
Moreover, our approach mainly distinguishes between objects of the same type via composition nodes.
Yet, we believe that there is a more effective strategy than merely assigning numbers to these objects.
A.2.3 Limitation concerning our experiments
Our experiments, which focus on CLIP training and retrieval tasks, demonstrate the benefits of our method and dataset from several perspectives.
Nonetheless, we believe this only represents a small part of what this new dataset and annotation method can offer.
Moving forward, we commit to exploring broader applications, particularly in text-to-image and image-to-text problems, to fully leverage our dataset’s potential.
A.3 Societal impact
Our paper introduces the GBC datasets and procedure, both aimed at advancing the development of multimodal models.
Specifically, the structured approach of GBC, designed to provide detailed descriptions, may help overcome representational biases inherent in existing captioning pipelines, offering more accurate descriptions of images.
The potential benefits of these advancements extend across a range of applications, such as assistive technologies and scientific research.
However, alongside these benefits, there are challenges including the potential spread of misinformation and concerns about privacy.
A comprehensive discussion of these broader societal impacts, both positive and negative, extends beyond the immediate focus of our methodological study.
Appendix B Dataset construction
In this appendix, we provide all the missing details about our dataset construction process that are not mentioned in Sections 3.2 and 3.3.
B.1 Query templates
To make the MLLM models fulfill the tasks described in Section 3.2, we perform COT prompting [66] with few-shot examples.
The four templates for our queries are shown in Figure 5 to 11.
We make the following remarks concerning the design of our prompts.
Prompt structure. We craft these prompts with the help of ChatGPT, which results in prompts that might be more complicated than necessary.
Meanwhile, we did notice that the inclusion of few-shot examples is crucial for the model to adhere to the required output formats.
Given that using always the same few-shot examples might significantly bias the model’s output, it could be beneficial to randomly retrieve examples from a diverse pool for each query, but we did not pursue this exploration.
[Single] and [Multiple] annotations. Since a detection model could output multiple candidate bounding boxes for an input text, we ask the MLLM to annotate each identified element with either [single] or [multiple].
We then proceed with slightly different algorithms in the two cases, to encourage the selection of either only one, or multiple bounding boxes.
In particular, we use respectively an NMS threshold of and for objects labeled with [single] and [multiple].
However, these labels do not necessarily dictate the final count of bounding boxes; multiple boxes may still be selected for items labeled [single], and vice versa.
Dynamically filled-in elements. To ensure that the response of the MLLM is relevant, the prompts are dynamic and reflect the content of the current image (the image query being the only exception).
Such information comes from previous queries and can be naturally retrieved for different queries.
The only nonobvious part is the hard coded hints for composition queries, which we explain below.
Hard coded hints for composition queries. After numerous attempts, we observe that LLaVA-1.6 struggles with accurately describing the composition of multiple objects in a scene, even when these objects are annotated with bounding boxes.
To overcome this limitation, we guide the models with hints generated programmatically using a set of predefined rules. Specifically, we begin by constructing a Euclidean minimum spanning tree based on the centers of the bounding boxes. We then select a random node as the root and perform a DFS (DFS) on the tree. During this search, we interleave descriptions of the edges, which detail the geometric relations between two objects based on the positions of their bounding boxes, with node descriptions. These node descriptions are added when an object is located at a particular extremity of the composition, such as the rightmost or top-left position.
B.2 Text classifiers
Both of our text classifiers are trained for binary classification using logistic loss. To determine whether a piece of text is suitable for object detection, we utilize a single linear layer added on top of the Jina Embedding.333https://huggingface.co/jinaai/jina-embeddings-v2-small-en Accessed: 2024-05-01 For the task of assessing whether two texts can represent the same object, we concatenate their Jina embeddings and process them through an MLP. This MLP includes layer normalization, a hidden layer that expands the input dimensionality by a factor of four, followed by SiLU activation and the final linear layer.
The dataset for the training of our text classifiers are prepared with the help of ChatGPT.
B.3 Other details for data annotation
We incorporate LLaVA-1.6 into our pipeline using llama.cpp.444https://github.com/ggerganov/llama.cpp Accessed: 2024-05-01
Moreover, to speed up the annotation process,
we utilize models quantized at different precision levels: the vision encoders at 6-bit precision, the LLM component of LLaVA-1.6 Mistral-7B at 5-bit precision, and the LLM component of LLaVA-1.6 Yi-34B at 3-bit precision.555
https://huggingface.co/cmp-nct/llava-1.6-gguf Accessed: 2024-05-01
We use the default hyperparameters for inference except for a temperature of 0.1 and context window of size of 5952 (note that LLaVA-1.6 can use up to 2880 image tokens).
We discard any responses that do not comply with our required format.
As for the object detection model, we use YOLO-Worldv2-X trained with input resolution of .666https://huggingface.co/wondervictor/YOLO-World/blob/main/yolo_world_v2_x_obj365v1_goldg_cc3mlite_pretrain-8698fbfa.pth Accessed: 2024-05-01
We set the confidence threshold to 0.05 and retain a maximum of six bounding boxes for each input text, selecting those with the highest confidence scores. We exclude any region whose size is smaller than 5,000. To prevent repetitive descriptions of the same element, we keep only those bounding boxes that occupy less than 80% of the current image region for detections arising from entity queries. Regarding node merging, we consider two bounding boxes to be overlapping if their intersection occupies more than 85% of the area of each bounding box involved.
B.4 Computation cost
We list below the major computation cost of our data preparation process.
GBC1M: With our processing pipeline, it takes an average of around 3 minutes to annotate each image on an A100 80G when all the queries are performed with LLaVA-1.6 Yi-34B.
As a result, annotating 1 million images took us around 6 days with 300 A100 80Gs.
GBC12M:
The average annotation time per image on an A100 80G is improved to 1 minute when relation and entity queries are performed with LLaVA-1.6 Mistral-7B.
This process is about twice slower on a V100 32G.
In this regard, our GBC12M dataset was compiled in roughly 6 days using 500 A100 80Gs and 1,000 V100 32Gs.
We also compute CLIP score for each caption using the DFN-5B model.777https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378 Accessed: 2024-05-01
This computation takes around 3 hours for every 10,000 images on a V100 32G.
Figure 5: The system prompt used for image query (first half).
Your task is to perform an in-depth analysis of a cropped image focusing on a requested object, like a "house". The process involves a step-by-step evaluation to identify the object’s presence, describe its features, craft concise captions, and assess any prominent objects.
Process Overview:
Verify Object Presence:
- Examine the image to determine if the specified object, or any instance of it, is present.
- State the presence with "Object Present: Yes" or "Object Present: No".
Provide Appropriate Caption (If Object Is Present):
- Provide a detailed description of the object, focusing solely on its features without reference to other elements in the image.
- The description should contain at most 50 words.
Assessment of Prominent Objects:
- Evaluate the described features to determine if any stand out for further description and are detectable by an object detection model. This is crucial for complex objects such as ’man’, ’woman’, ’family’, ’couple’, ’cat’, or ’house’, where components or distinctive attributes are significant. For example, when analyzing ’woman’, consider elements like dress [single], skirt [single], or hair [single] as prominent features. For simpler objects like ’cup’ or ’chair’, additional descriptions may not be needed.
Identification of Prominent Features (If Applicable):
- If there are prominent features identified, list and format these features for potential detection by an object detection model.
- Ensure these features are parts or components of the main object and not the entire object itself.
- Use [single] for unique, standalone items, and [multiple] for features present more than once, such as roof [single] or windows [multiple].
- Group similar items under a single [multiple] label rather than describing them separately, even if individual descriptions were provided in the detailed caption. For example, multiple distinct windows in a house should be labeled as windows [multiple] rather than individually enumerated.
- For groups like families or couples, identify members separately (e.g., man [single], woman [single]) rather than as a collective unit. This contrasts with grouping similar inanimate objects (e.g., windows [multiple]), where individual distinction isn’t necessary.
- Consistency with the caption: Ensure that the features identified as [single] or [multiple] are also mentioned in the caption.
Example Responses:
Example 1: Object Not Present
Object Presence: No
Detailed Caption: N/A
Prominent Features: N/A
Identification of Prominent Features: N/A
Example 2: Object Present Without Prominent Features (requested object: "cup")
Object Presence: Yes
Detailed Caption: A simple ceramic cup on a wooden table. The cup has a smooth, unadorned surface and a standard curved handle on one side.
Prominent Features: No
Identification of Prominent Features: N/A
Example 3: Object Present With Prominent Features (requested object: "family")
Object Presence: Yes
Detailed Caption: A family of four is captured enjoying a sunny day in the park. The father, in casual attire, is engrossed in setting up a picnic spot, while the mother, donned in a summer dress, is laying out a feast on a blanket. Nearby, two children, a boy and a girl, engage in playful antics; the boy is kicking a football with fervor, and the girl, adorned in a light frock, is gleefully chasing bubbles.
Prominent Features: Yes
Identification of Prominent Features:
- Father: [single]
- Mother: [single]
- Boy: [single]
- Girl: [single]
Example 4: Object Present With Prominent Features (requested object: "car")
Object Presence: Yes
Detailed Caption: A vintage car in pristine condition, with shiny chrome bumpers and classic spoke wheels. The car’s body is painted in a vibrant red, and the leather interior is visible through the clear windows. A unique hood ornament adorns the front, adding to the car’s elegance.
Prominent Features: Yes
Identification of Prominent Features:
- Chrome bumpers: [single]
- Wheels: [multiple]
- Hood ornament: [single]
boat
boat
Your role is to analyze images containing objects within pre-labeled bounding boxes and describe the compositional arrangement of these objects based on provided hints. You will then provide general descriptions that apply to all the objects collectively.
Input Image Format Explanation:
- The image will feature objects of interest, each enclosed within a bounding box.
- Each bounding box will be numbered centrally to uniquely identify it.
- The objects will be similar in nature (e.g., all dogs) and positioned within a scene.
Utilizing Hints for Analyzing Composition:
- Begin by reviewing the hints provided regarding the spatial arrangement of the objects.
- These hints may specify the relative positions of objects (e.g., "Object 3 is in the top right corner").
- Use the hints to guide your description of how the objects relate to each other within their bounding boxes.
Output Format:
- Composition Description: Start with "Composition:" followed by a description informed by the hints and using the bounding box numbers. This description should elucidate the spatial arrangement of the objects as per the hints.
- General Descriptions: Provide observations that apply to all objects within the specified group, excluding unrelated elements or background details. Preface this section with "General descriptions:".
Additional Guidelines:
- Describe the spatial arrangement of objects without inferring spatial relations from the sequence of numbers.
- Utilize clear spatial language to articulate the composition.
- The description should reflect the actual visual composition, not the order of numbers in the bounding boxes.
Examples:
Example for 3 Dogs in Bounding Boxes:
Query Prompt: "Please describe the composition of the 3 dogs in the bounding boxes, followed by some general descriptions that apply to all dogs."
System Response:
Composition: Dog 3 is in front, with dog 2 to the left and dog 1 to the right.
General descriptions:
- The three dogs are aligned in a row on the grass.
- They share similar sizes and features, suggesting they may be from the same breed.
Additional Examples:
For 5 Flowers in a Garden Bed in Bounding Boxes:
Composition: Flower 4 takes a central position, flanked by flower 2 and flower 3 on either side, while flower 1 and flower 5 bookend the arrangement at the outer edges.
General descriptions:
- Each flower is in full bloom, indicating a peak growing season.
For 2 Cats in a Window in Bounding Boxes:
Composition: Cat 1 is positioned on the left side of the window sill, while cat 2 is curled up on the right.
General descriptions:
- Both cats are basking in the sunlight coming through the window.
- Their relaxed postures suggest a shared sense of comfort and tranquility.
- boat 2 is to the right of boat 3
- boat 3 is to the right of boat 1
- boat 1 is on the left side of the composition
Your role involves analyzing the spatial and direct interactions between pre-identified elements within an image, described through annotations like [beach], [turquoise waters], [people], [shoreline], [lush vegetation]. Your task is to objectively describe how these elements are related or positioned relative to each other within the scene.
Steps for Identifying Objective Relations:
1. Review Annotated Elements: Start by examining the list of annotated elements. Understand the nature of each element as it is presented in the image.
2. Identify Spatial Positions: Determine the spatial positioning of these elements in relation to each other. Focus on direct relationships such as touching, overlapping, or proximity without implying any further interpretation.
3. Describe Direct Interactions: Look for and describe any direct interactions between elements, such as one element supporting another, blocking, or leading into another.
4. Format Relations List: Provide your findings as a list of bullet points. Each bullet should detail a direct and observable relationship between two or more elements, using their annotated identifiers for reference.
Example Relations Based on Annotated Elements:
For elements: [beach], [turquoise waters], [people], [shoreline], [lush vegetation], you might reply:
- The [people] are standing on the [beach], with the [lush vegetation] to their left.
- [Turquoise waters] lap against the [beach] at the [shoreline], with [people] scattered along its edge.
- [Lush vegetation] flanks the left side of the [beach], providing a natural border.
- The [shoreline] separates the [beach] from the [turquoise waters].
- To the right of the [lush vegetation], the [beach] stretches towards the [turquoise waters].
For another set of elements: [eagle], [snake], [wings], you might reply:
- The [eagle] has its [wings] spread above the [snake].
- The [snake] is positioned below the [eagle].
- The [eagle]’s claws are near or in contact with the [snake].
Guidelines for Reporting Relations:
1. Ensure descriptions are based solely on visible or directly inferable relationships without adding interpretations or assumptions.
2. Maintain clarity and precision in articulating the spatial and interactional dynamics between elements.
3. Stick to objective descriptions that reflect the physical and observable aspects of the elements’ relationships.
4. Only answer the list of bullet points without anything before or after.
5. Do not include any bullet point with 1 or even 0 elements.
- Visible Relationships Only: Report relationships that are clearly depicted in the image. If no clear relationships are visible, state "No visible relationships."
- Objective Descriptions: Keep descriptions factual and based solely on what can be seen in the image.
- Avoid Assumptions: Do not infer or assume any relationships that aren’t clearly shown in the image.
- Bullet Point Format: Present each observable relationship as a separate bullet point, avoiding any descriptive text not related to the direct relationships.
- No Relation Inference: Refrain from implying relationships or positions that are not explicitly shown. If elements are simply present without any discernible interaction, it is acceptable to say "Elements are present without visible interaction."
- Avoid Single Element Points: Do not include bullet points that mention only one element or have no elements at all. Each bullet point must reference the relationship between two or more elements.
4
eiffel tower, sky, trees, boat, water
Appendix C Dataset information
In this appendix, we provide information about dataset release,
dataset statistics, and visualizations of a few examples from our GBC10M dataset.
C.1 Data release and licensing
Our datasets are available at https://huggingface.co/graph-based-captions, released under the Apple Sample Code License.888Should we receive the necessary approvals, we may transition to a less restrictive license in the future.
Following CC12M, we include URLs to images along with captions generated through our GBC procedure, all stored in JSON lines format.
Comprehensive documentation including a dataset card and croissant metadata is provided in the data repository.
We are committed to maintaining the dataset to ensure its long-term public accessibility.
Personal identifiable information and offensive content.
Our dataset comprises only captions generated by MLLM models (LLaVA 1.6 Yi-34B and LLaVA 1.6 Mistral-7B), which were trained on carefully curated data.
The images, sourced from CC12M, are generally free from offensive content.
In particular, CC12M is the result of a filtering operation involving adult content detection on images and their captions.
While CC12M images may include human faces, we do not host the images directly; only the URLs are provided.
Additionally, we conduct toxicity check with Detoxify [25] on a subset of examples in GBC dataset and find no harmful contents.
While it was not possible to manually examine all the samples produced by GBC pipeline, we believe that the protective measures of the source dataset and model are sufficient to avoid both harmful content, and privacy leakages.
Author statement.
We, the authors, accept full responsibility for any violation of rights.
C.2 Dataset statistics
In this section, we provide statistical insights into the GBC1M and GBC10M datasets. In particular, we zoom in on the statistics at image, vertex, edge, and caption levels, and present distributions of several key metrics including for example caption length, region size, and CLIP score.
Since most of these metrics exhibit long-tailed distributions, we often group excessively large values into a single histogram bin for better visualization.
C.2.1 Image and graph statistics
C.2.2 Vertex statistics
Dataset
# Images
# Regions / Image
COCO [41]
123,000
7
Visual Genome [33]
108, 249
42
Objects365 [55]
638,000
16
Open Images [34]
1.7M
8
BigDetection [5]
3.5M
10
SA-1B [32]
11M
100
AS-1B [65]
11M
110
DCI [14]
7,805
40
GBC1M (ours)
1.1M
11
GBC10M (ours)
10.1M
11
Table 6: Comparison of number of regions per image among several vision-language datasets with region-based annotations.
We use the statistics reported in the original paper although some datasets, such as COCO and Open Images have been updated after their initial release.
Moreover, for Open Images we report the number for the training set with bounding box annotation [34, Tab. 5].
For DCI, we compute the average number of regions per image ourselves using their open-sourced dataset with 7,805 images as this number is not reported in [14].
We have shown previously that our datasets contain an average of 12 vertices per graph.
This translates to 11 regions per image after excluding the root node that represents the entire image.
We compare this number with several other vision-language datasets with region-based annotations in Table 6.
As one can see, this number aligns well with many of these datasets, particularly those used for detection, such as COCO and Object365.
However, it lags behind compared to Visual Genome and more recent datasets with dense annotations, such as AS-1B and DCI.
We believe this discrepancy can be attributed to both the top-down design of our annotation process, which tends to overlook less significant components of the images, and the limitations of the detection model used.
Notably, both AS-1B and DCI utilize Segment Anything [32] to identify regions of interest. Segment Anything is trained on SA-1B, which has much denser annotations compared to the object detection datasets used for training Yolo-World.
C.2.3 Edge statistics
(a) GBC1M
(b) GBC10M
(a) GBC1M
(b) GBC10M
C.2.4 Caption statistics
C.3 Examples from GBC10M
As a complement to the dataset statistics presented in the previous section, we showcase a few illustrative examples from GBC10M in Figures 30 and 31.
These examples demonstrate the varying levels of graph complexity across our dataset.
The number of nodes varies from just a few (first example in Figure 30) to over 10 (third example in Figure 30 and the example in Figure 31).
In most cases, this complexity aligns with the visual complexity of the corresponding image.
On the other hand, these examples also reveal limitations arising from the object detection models used.
For instance, in the Messe example from Figure 30, the detection model incorrectly identifies a standing priest as a “kneeling figure”.
Similarly, in Figure 31, two of the three nodes labeled “trunk” are derived from tree nodes and erroneously associated with the elephant’s trunk or other non-trunk objects on the elephant.
These limitations become particularly severe in the Regalia example of Figure 30, where the presence of more specific objects like crowns, scepters, bracelets, and earrings leads to frequent confusion by the object detection model.
Next, we focus on the captions associated with these images.
A subset of these captions is presented in Sections C.3 and C.3.
We observe that hallucination is particularly important for detailed captions.
These erroneous descriptions can then be inherited by the shorter captions derived from them.
We also note there are situations where the model describes an object that actually does not exist in the corresponding region of the image, such as the caption for “scepter 1” in the Regalia example.
As we can see from the figure, in the corresponding bounding box, there is no scepter visible but only a crown on a wooden base.
In spite of these inaccuracies in object detection and captioning, the overall graph structure and captions still align well with the images.
On top of this, the granularity of our descriptions significantly enhances the descriptive power of our dataset, allowing for a more nuanced understanding of the visual content.
Messe [Figure 30]. A priest holds a chalice aloft while another figure kneeling figure kneels on the floor, set against a backdrop of architectural details and ornamentation within what appears to be a religious setting.
Regalia [Figure 30].
Two people stand behind a display case containing a crown, scepter with a blue gem, and a golden orb with a red gem, all under natural light from a window or glass panel within an indoor setting.
Elephant [Figure 31].
A man rides atop a bench strapped on a elephant drinking from a riverbank, surrounded by trees under a clear sky.
Detailed
Captions
Flame [Figure 30].
The image captures a close-up view of a blue flame emanating from a small metal object, which appears to be a lighter or torch. The flame has a vibrant yellow hue at its base, transitioning to a bright blue at its peak. The flame’s shape is irregular with wisps extending outward from its core, suggesting it’s in motion or has been recently ignited. The metal object has a cylindrical shape with a pointed tip from which the flame emerges. The background is dark, providing a stark contrast that accentuates the flame’s colors and form.
Messe [Figure 30]. The image portrays a religious scene set within what appears to be a church or chapel. At the center of the composition stands a priest, dressed in traditional religious attire with a red robe and a white cowl. He holds a golden chalice aloft with both hands, suggesting he may be performing a sacrament or ritual. To his right, another figure, possibly another priest or religious figure, kneels on the floor, seemingly in prayer or reverence. The background features ornate architectural details, including arches and intricate patterns on the walls, indicative of Gothic or similar architectural styles. The overall atmosphere suggests a solemn or sacred moment within a religious ceremony or service.
Regalia [Figure 30].
The image captures a scene where two individuals are standing behind a display case containing various items. The display case houses a collection of ornate jewelry pieces, including a crown with intricate detailing, a scepter with a blue gem at its top, and a golden orb with a red gem.
The individuals are dressed in pink shirts and are positioned behind the display case, which has a reflective surface. The background suggests an indoor setting with a window or glass panel allowing natural light to illuminate the scene.
Elephant [Figure 31].
The image captures a serene scene at a riverbank where a man is riding on the back of a large elephant. The elephant, with its majestic gray skin, is partially submerged in the water, drinking from it. The man, dressed in casual attire, sits comfortably on a wooden bench strapped securely on the elephant’s back. The bench is adorned with colorful cushions for added comfort during the ride.
The backdrop features lush greenery with trees lining the riverbank, adding to the tranquil atmosphere of the scene. The overall setting suggests a peaceful interaction between humans and nature.
Messe—priest [Figure 30].
A priest dressed in traditional religious attire, which includes a red robe with gold trim, a white cassock, and a red sash. He holds a cross in his right hand.
Regalia—scepter 1 [Figure 30].
A gold scepter with a cross atop it. The scepter has intricate designs and patterns throughout its length.
Elephant—elephant [Figure 31].
A large elephant with a long trunk is seen walking through a body of water. Its skin appears rough and wrinkled, typical of elephants. The elephant has small tusks and large ears that are characteristic of this species.
Relation
Captions
Flame—flame/metal object [Figure 30]. The flame is positioned above the metal object.
Messe–priest/kneeling figure [Figure 30]. The priest is standing in front of the kneeling figure.
Regalia—crown/scepter [Figure 30]. The scepter is positioned next to the crown.
Elephant—elephant/riverbank/trees [Figure 31].
The elephant is standing near the riverbank with trees in the background.
Composition
Captions
Messe–kneeling figure [Figure 30]. Kneeling figure 1, positioned at the top right, appears to be in a state of prayer or reverence, while kneeling figure 2, located at the bottom left, seems to be in a similar posture, possibly indicating a shared moment of devotion or reflection.
Regalia—scepter [Figure 30].
Scepter 2, which is in the bottom left corner, has a golden handle with a blue gemstone at its center, while scepter 1, positioned above scepter 2, features a golden handle with a red gemstone at its center. Both scepters are ornate, with intricate designs and a regal appearance.
Elephant—riverbank [Figure 31].
Riverbank 1 is positioned above Riverbank 2, with Riverbank 2 located at the bottom of the composition.
Multi-Entity
Captions
Messe–kneeling figure [Figure 30]. Both figures are depicted in a posture commonly associated with prayer or worship, suggesting a religious or spiritual context for their actions.
Regalia—scepter [Figure 30].
The gemstones in their handles add a touch of elegance and value to each scepter.
Elephant—riverbank [Figure 31].
They are situated near a body of water, which suggests a peaceful, natural setting.
Appendix D Algorithm details
This appendix provides missing details about our architecture and training objective.
D.1 Structure-aware cross attention
We define below mathematically the SACA layer.
For this, we denote by the children of caption in caption graph
and write the features of in the input of our SACA layer as .
Recall also that with represents the set of token positions in the source caption that we map the edge label to.
Then, the SACA layer maps each feature vector to
where implements the standard multi-head attention mechanism.
Note that we average across the results from all the relevant captions that describe this token, as we show in Figure 4.
As a side note, we highlight that with SAHA, information is only propagated from each node to its direct parent within a block.
Consequently, the number of blocks must exceed the depth of the to ensure that information reaches the root node from all levels of the graph.
D.2 Multi-positive contrastive loss
To pair multiple positive captions to an image, we extend standard contrastive loss [50] into multiple-positive contrastive loss, as also considered in prior studies [14, 16]. Specifically, consider a batch of images , where each image is associated with captions , we utilize the following loss function to account for multiple positive texts per image:
where , is a learnable temperature parameter, and is a normalizer. On the other hand, each caption still only has one paired image. Therefore, we use the standard contrastive loss on for text-to-image alignment:
Appendix E Experimental details
This appendix presents further details about our experiments that are omitted in Section 5.
E.1 Data filtering
For the computation of CLIP score, we split any caption that contains more than 77 tokens into individual sentences, compute the score for each of these sentences, and compute the average of these scores.
Then, we start by filtering out images whose short synthetic captions have CLIP scores that are lower than the 5% quantile.
After this, we consider three filtering strategies depending on the annotation formats.
Long caption. In this case, we just further filter out a portion of original captions and long captions with the lowest CLIP scores (by considering the 5% quantiles from the non-filtered dataset).
GBC. Naive CLIP filtering and tokenizer truncation could break the graph structure as some of the edge labels would not appear in the captions of its source node anymore after these operations.
We address this issue by filtering out the captions following the reverse of a topological ordering of the graph, drop a node along with its in edges when all its captions and children get filtered, and otherwise, if necessary, add bag-of-words captions that collects edge labels from the remaining out edges of a node to ensure all these labels still appear in some captions of this node.
Moreover, we split the captions whose length are longer than 77 tokens into concatenations of sentences that fit within this limit, and drop any caption which contains sentences that are of more than 77 tokens.
Short and Region. We follow the strategy mentioned in GBC, but use selected types of captions.
Moreover, bag-of-words captions are not used.
We remark that the filtering procedure is only applied to the training set, and not the GBC test set.
E.2 Dynamic batch size
Given the varying sizes of our graph, setting a fixed number of images per batch could result in out-of-memory errors unless we opt for a conservatively small batch size.
To overcome this challenge, we implement a dynamic batching strategy for the setups where the number of captions per image is in principle unbounded.
This encompasses notably region, GBC-captions, and GBC-graph.
With this strategy, we ensure that the number of captions, and, in the case of GBC-graph, the number of edges, that are included in each batch do not exceed a certain limit.
In this regard, the batch size that we report in Section 5 is actually just an upper bound on the number of images included in each batch.
More specifically, we set this limit based on the number of average captions/edges per image in the filtered dataset.
For example, for GBC-captions and GBC-graph we have in average 17.61 captions per graph.
We thus set the limit on caption number to on each GPU (as mentioned in Appx. E.4, we use 64 GPUs for most of our experiments, which gives a batch size of per GPU).
E.3 Hyperparameters for CLIP training
We used a consistent set of hyperparameters for all model training runs, as detailed in Table 11.
The sole exception is training with original CC12M captions, where we used a larger batch size of 8,192 to ensure the model sees a comparable number of texts as during training with both short synthetic and original captions.
For this specific setup with the larger batch size, we reported evaluation results from the EMA checkpoint at the end of epoch 15, for it achieving the best performance among the evaluated checkpoints.
For GBC-graph, we drop the edges with probability so that the model also learns how to match images with short captions.
E.4 Computation cost
We train all CLIP models on A100-80G GPUs. As training with different annotation formats requires varying size of GPU memory, we use different total numbers of GPUs to ensure the same batch size. Specifically, we utilize 16 GPUs for training with Short captions, and utilize 64 GPUs for training with all other annotation formats. We list the corresponding time required for training with different annotation formats in Table E.4.
Training time (hr) 22.7 19.7 29.9 38.9 20.3 43.0
E.5 Evaluation details
Our evaluation uses the validation set of ImageNet-1k [51] and the test sets of Flickr30k [47] and MS-COCO [41].
For SugarCrepe [27] we report the average performance across all variants.
As for ShareGPT4V [76], we use a subset of size from ShareGPT4V-cap100k.
These images were also used for LLaVA training.999https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain Accessed: 2024-05-14
When each image is paired with multiple captions, we only select one of them.
The evaluation setups with ADE20K and DCI are more involved, as we explain below.
Evaluation on ADE20K. We evaluate the quality of CLIP models’ image encoder for dense prediction tasks like image segmentation by performing full finetuning on ADE20k [76] dataset. We follow the same setup as described in [62, 63] where we use a ViTDet style feature pyramid network with UperNet [67] head. All models were trained using the MMSegmentation library [11]. We sweep through peak learning rate for all the results reported in the paper and the ranges are listed in Table 11.
Evaluation on DCI. We perform text-to-image and image-to-text evaluations on DCI [61] using either long captions or concatenated captions.
The long captions are marked as extra_caption in the released DCI dataset.
We filter out samples with empty long captions, resulting in a subset of 7,602 images for evaluation with long captions.
Regarding evaluation with concatenated captions, we leverage the full set of 7,805 images.
We retain masks containing summary captions (these are masks with bounding boxes larger than ).
If the human-annotated caption contains fewer than 77 tokens and is longer than the first summary caption, we use it.
Otherwise, we use the first summary caption.
For concatenation, we follow the BFS (BFS) order based on the provided tree structure between the masks.
Appendix F Additional results and experiments
In this appendix, we present additional ablations that we have performed but were not presented in the main paper due to space constraints.
Epoch
Batch size
# Tokens
ImageNet
Flickr
COCO
Share-
GPT4V
DCI-
concat
GBC
test
Short
10 4,096 77
38.8 64.8 38.7 79.1 57.5 87.8
512
39.0 64.7 39.3 86.7 56.4 89.7
10
33.2 59.1 34.7 74.0 52.3 83.4
28 16,384 77
40.0 67.4 38.7 80.6 58.1 88.6
40
39.0 65.7 37.3 79.9 57.8 88.6
GBC-captions
40.8 70.0 43.0 80.4 64.1 91.2
GBC-concat
10 4,096 77
39.0 66.1 40.0 90.5 69.9 94.8
GBC-graph
38.2 67.1 40.8 77.1 61.5 95.9
F.1 Matching compute resource for training with short captions
All our models presented in Section 5 used 8 nodes for training, except for the models trained on short captions, which only used 2 nodes.
This raises the question of whether the performance gap could be bridged by providing more computational resources to this setup.
To address this, we specifically considered two modifications that would naturally necessitate using more nodes for training with short captions: (i) extending the context length to 512, as done for training with Long and GBC-concat captions, and (ii) using a batch size that is four times larger, i.e., a batch size of 16,384 instead of 4,096.
All other hyperparameters remained unchanged.
We then trained the models on 8 nodes, each with 8 GPUs, as in the other setups, which resulted in training times of 18 and 48 hours for the two modifications respectively.
The evaluation results are presented in Appendix F.
Training with extended context length.
Provided that the models are only trained with short captions, we do not expect any tangible benefit from extending the context length.
Yet, surprisingly, while this is indeed the case for classic benchmarks such as ImageNet, Flickr, and COCO, we do observe a significant performance boost on ShareGPT4V retrieval, suggesting that the longer context length is still beneficial for retrieval with long caption even though the model is not explicitly trained for this task.
On the other hand, we do not observe any benefit when evaluated using concatenated caption from DCI.
Finally, we also get a slight performance improvement on GBC test, and it turns out this improved performance is achieved by performing retrieval using the long caption.
This is in line with the performance gain that we observe for the ShareGPT4V benchmark.
Training with larger batch size.
More interestingly, CLIP is known to perform better when trained with a large batch size, so
we might be able to bridge the performance gap by simply including more images and captions in each batch.
To enable a fair comparison for this setup, we report evaluation results from three checkpoints at varying training stages in Appendix F. These checkpoints are chosen to align with key training milestones.
Number of images seen: We consider the EMA checkpoint at the end of epoch 10 to align the number of images seen.
Number of iterations: We include the EMA checkpoint at the end of epoch 40 to compare models at a fixed number of training iterations.
Best performing checkpoint: Additionally, we report results for the EMA checkpoint at the end of epoch 28, as it gives the best performance among all evaluated checkpoints (see Figure 32).
As we can see from the table, while the use of a larger batch size indeed leads to better performance on ImageNet and Flickr, the results still lag behind those achieved with GBC-captions. This discrepancy underscores the importance of including multiple captions per image to enhance performance.
F.2 The importance of multi-positive contrastive loss
Short
38.8 35.2
64.8 61.0 38.7 36.7 76.0 74.4 -2.75
Long
39.6 30.5
65.8 56.8 40.1 33.5 77.0 74.0 -6.93
GBC-captions
40.8 31.9
70.0 58.3 43.0 33.2 76.7 73.3 -8.45
F.3 Impact of caption type on CLIP training
T2I
I2T
T2I
I2T
T2I
I2T
Short
56.3 73.2 30.7 46.7
57.5 57.5
38.8 76.0 42.0
Region
58.3 76.6 31.5 49.1
61.8 61.5
38.5 75.6 43.5
GBC-Relation
60.4 76.5 34.8 52.5
62.0 61.4
41.5 76.4 44.5
GBC-captions 60.6 79.3 34.1 51.9
64.1 63.4
40.8 76.7 45.0
F.4 Impact of the underlying graph on retrieval
T2I
I2T
Groundtruth
Last token
Random token
T2I
I2T
T2I
I2T
T2I
I2T
T2I
I2T
T2I
I2T
GBC-graph 84.3 85.2
95.7 96.1
91.7 92.4
95.1 95.4
94.8 94.9
91.8 92.5
F.5 Evaluating at non-EMA checkpoints
CC12M 37.1 50.5 27.9 39.4 47.2 49.5
Short 37.5 62.0 36.1 74.5 77.3 86.9
Long 38.5 64.5 38.4 75.8 93.5 95.5
Region 40.3 68.6 40.8 76.1 78.9 91.8
GBC-captions
41.3
70.6
43.1
76.4
80.1 91.9
GBC-concat
38.2 63.4 37.1 74.9 89.8 96.1
GBC-graph
39.9 68.5
40.9 74.6 77.5 96.2
Appendix G Image attributions
All the images that we show in this paper come from Wikimedia Commons.
We provide in Table 18 the exact source urls and license for each of the images.
The urls to the CC BY-SA 3.0 and GFDL 1.2 licenses are respectively https://creativecommons.org/licenses/by-sa/3.0/ and https://www.gnu.org/licenses/old-licenses/fdl-1.2.txt.
Image
Source URL
License
Figure 1
https://commons.wikimedia.org/wiki/File:Tartu_raudteejaama_veetorn,_2010.JPG
CC BY-SA 3.0
Figure 2
https://commons.wikimedia.org/wiki/File:Eiffel_Tower_from_north_Avenue_de_New_York,_Aug_2010.jpg
CC BY-SA 3.0
Figure 30
Flame
https://commons.wikimedia.org/wiki/File:Flametest--Na.swn.jpg
CC BY-SA 3.0
Messe
https://commons.wikimedia.org/wiki/File:Messe_mit_Wandlungskerze_Beuron.jpg
Public domain
Regalia
https://commons.wikimedia.org/wiki/File:Crown,_sceptre,_orb_%26_key_of_the_King_of_Sweden_2014.jpg
Public domain
Figure 31
https://commons.wikimedia.org/wiki/File:Indian-Elephant-444.jpg
GFDL 1.2
Table 18: Source URLs and licenses of the images shown in this paper.